# Off-policy Evaluation with Deeply-abstracted States

### v2

@@ D.2 Proof of Lemma 1
    $Q_{j,\phi}^{\pi}(a,x)=\sum_{t=1}^{j}\gamma^{t-1}\mathbb{E}^{\pi}\left[R_{t}|\phi(S_{1})=x,A_{1}=a\right].$ (D.5)
    Similarly, we can show that
    $=$ $\mathbb{E}\left[(1-\gamma)^{-1}w_{\phi}^{\pi}(A,\phi(S))\underbrace{\mathcal{R}_{\phi}\big{(}A,\phi(S)\big{)}}_{\mbox{reward-irrelevant}}\right]$
    $=$ $\sum_{t=1}^{+\infty}\gamma^{t-1}\mathbb{E}^{\pi}({R}_{t})$
    $=$ $\mathbb{E}[f_{3}(w^{\pi})].$
    $\rho_{j,\phi}^{\pi}=\frac{\pi_{\phi}(A_{j}|\phi(S_{j}))}{b_{j,\phi}(A_{j}|\phi(S_{j}),A_{j-1},\phi(S_{j-1}),\ldots,A_{1},\phi(S_{1}))}.$ (D.6)
    Under $\pi$-irrelevance, the numerator is well-defined and identifiable. However, unlike the standard IS ratio where the denominator depends solely on the current state, the denominator in (D.6) depends on the entire history. Notice that the denominator is essentially the data distribution of $A_{j}$ given $\phi(S_{j}),A_{j-1},\phi(S_{j-1}),\ldots,A_{1},\phi(S_{1})$, it is thus identifiable from the offline data as well.
    $\mathbb{E}(\rho_{1:t}^{\pi}R_{t})=\mathbb{E}(\rho_{1:t,\phi}^{\pi}R_{t}),$ (D.7)
    Additionally, since the generation $A_{t}$ depends only on $S_{t}$, the inner expectation equals $\mathbb{E}(\rho_{1:t-1}^{\pi}|S_{t})$ which can be further shown to equal to $p_{t}^{\pi}(S_{t})/p_{\infty}(S_{t})$. This allows us to represent the left-hand-side of (D.7) by
    $\mathbb{E}\Big{[}\frac{p_{t}^{\pi}(S_{t})}{p_{\infty}(S_{t})}\rho^{\pi}(A_{t},S_{t})R_{t}\Big{]}.$ (D.8)
    Using similar arguments to the proof of Fisher consistency of MIS estimator, under reward-irrelevance, (D.8) can be shown to equal to
D.3 Proof of Theorem 1
We establish the Fisher consistencies of SIS, MIS, Q-function-based method and DRL one by one.
  • Fisher consistency of SIS. Notice that $\rho^{\pi}$-irrelevance directly follows from the definition of backward-model-irrelevance and $\pi$-irrelevance. It follows from Lemma D.1 that $\mathbb{E}[f_{2}(\rho^{\pi})]=\mathbb{E}[f_{2}(\rho^{\pi}_{\phi})]$.
    $\begin{split}&\mathbb{E}[w^{\pi}_{t}(A_{t},S_{t})\rho^{\pi}(A_{t+1},S_{t+1})|A_{t+1}=a,S_{t+1}=s]\\ =&\sum_{a^{\prime},s^{\prime}}w^{\pi}_{t}(a^{\prime},s^{\prime})\frac{\pi(a|s)}{b(a|s)}\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime}|A_{t+1}=a,S_{t+1}=s)\\ =&\sum_{a^{\prime},s^{\prime}}\frac{\mathbb{P}^{\pi}(A_{t}=a^{\prime},S_{t}=s^{\prime})}{\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime})}\frac{\pi(a|s)}{b(a|s)}\frac{\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime},A_{t+1}=a,S_{t+1}=s)}{\mathbb{P}(A_{t+1}=a,S_{t+1}=s)}\\ =&\sum_{a^{\prime},s^{\prime}}\mathbb{P}^{\pi}(A_{t}=a^{\prime},S_{t}=s^{\prime})\frac{\pi(a|s)}{b(a|s)}\times\frac{\mathbb{P}(A_{t+1}=a|S_{t+1}=s,A_{t}=a^{\prime},S_{t}=s^{\prime})}{\mathbb{P}(A_{t+1}=a,S_{t+1}=s)}\mathbb{P}(S_{t+1}=s|A_{t}=a^{\prime},S_{t}=s)\\ =&\sum_{a^{\prime},s^{\prime}}\mathbb{P}^{\pi}(A_{t}=a^{\prime},S_{t}=s^{\prime})\frac{\pi(a|s)}{\mathbb{P}(A_{t+1}=a,S_{t+1}=s)}\mathbb{P}(S_{t+1}=s|A_{t}=a^{\prime},S_{t}=s)\\ =&\frac{\mathbb{P}^{\pi}(A_{t+1}=a,S_{t+1}=s)}{\mathbb{P}(A_{t+1}=a,S_{t+1}=s)}\\ =&w^{\pi}_{t+1}(a,s),\end{split}$ (D.9)
    $\begin{split}&\sum_{s^{\prime}\in\phi^{-1}(x^{\prime})}\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime}|A_{t+1}=a,S_{t+1}=s^{(1)})\\ =&\sum_{s^{\prime}\in\phi^{-1}(x^{\prime})}\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime}|A_{t+1}=a,S_{t+1}=s^{(2)}),\end{split}$ (D.10)
    whenever $\phi(s^{(1)})=\phi(s^{(2)})$.
    Indeed, by (5), we obtain that
    Consequently,
    $w^{\pi}_{t+1}(a,s^{(1)})$
    $=$ $w^{\pi}_{t+1}(a,s^{(2)}),$
    where the second last equation follows from $w_{t}^{\pi}$-irrelevance, $\rho^{\pi}$-irrelevance and (D.10). This yields $w_{t+1}^{\pi}$-irrelevance, and subsequently $w^{\pi}$-irrelevance, by induction. By Lemma D.1, $w^{\pi}$-irrelevance further yields $\mathbb{E}[f_{3}(w^{\pi})]=\mathbb{E}[f_{3}(w^{\pi}_{\phi})]$.
    Based on the aforementioned arguments, we can show that
    where the last term on the RHS is well-defined according to (D.10). Suppose we have shown the identifiability of $w_{t,\phi}^{\pi}$. Then each term on the RHS is identifiable. This proves the identifiability of $w_{t+1,\phi}^{\pi}$. By induction, $w_{t,\phi}^{\pi}$ is identifiable for each $t\geq 1$.
    Using the change of measure theorem, the numerator equals
    which is identiable from the data.
    Now, the identifiability can be readily obtained if we show $(\phi(S_{t}),A_{t},R_{t})_{t\geq 1}$ remains an MDP. In that case, standard Q-learning algorithms can be applied to such a reduced MDP to consistently identify $Q_{\phi}^{\pi}$. Such an MDP property will be proven in Section D.4.2 under a more challenging setting that allows the behavior policy to be history-dependent.
  • Fisher consistency of DRL. Due to the double robustness property of DRL, the conclusion directly follows from the last conclusion of Theorem D.1 and the first two conclusions of Theorem 1.

### v3

@@ D.1 Proof of Theorem 1
    $=$ $\mathbb{P}_{n}\left[\gamma\widehat{\alpha}^{\top}\varphi(A,S)\sum_{a^{\prime}}\pi(a^{\prime}|S^{\prime})\varphi\left(a^{\prime},S^{\prime}\right)^{\top}\beta-\widehat{\alpha}^{\top}\varphi(A,S)\varphi(A,S)^{\top}\beta\right]$
    $+(1-\gamma)\frac{1}{N}\sum_{i=1}^{N}\left[\sum_{a}\pi(a|S_{i,1})\varphi\left(a,S_{i,1}\right)\right]\widehat{\beta}$
    $=$ $(1-\gamma)\frac{1}{N}\sum_{i=1}^{N}\sum_{a}\pi(a|S_{i,1})\widehat{Q}^{\pi}\left(a,S_{i,1}\right)+\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}\widehat{w}^{\pi}\left(A_{i,j},S_{i,j}\right)$
    ×$\times\left(\gamma\sum_{a}\pi(a|S_{i,j+1})\widehat{Q}^{\pi}\left(a,S_{i,j+1}\right)-\widehat{Q}^{\pi}\left(A_{i,j},S_{i,j}\right)\right).$
    This means $\widehat{J}_{\mathrm{MIS}}(\widehat{w}^{\pi})=\widehat{J}_{\mathrm{DLR}}(\widehat{Q}^{\pi},\widehat{w}^{\pi})$. Thus, we have
    $\mathrm{MSE}_{\mathrm{MIS}}=\mathrm{MSE}_{\mathrm{DRL}}.$
  • MSE of SIS. Our SIS estimator based on the dataset $\mathcal{D}$ is
    $\widehat{J}_{\mathrm{SIS}}\left({\rho^{\pi}}\right)=$ $\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}\gamma^{j-1}\rho_{1:j}^{\pi}(A_{i,j},S_{i,j})R_{i,j},$
    where $T$ is the termination time. Direct calculations yield
    $\mathrm{MSE}_{\mathrm{SIS}}=$ $\mathbb{E}[\widehat{J}_{\mathrm{SIS}}\left({\rho^{\pi}}\right)-J(\pi)]^{2}$
    $=$ $\mathbb{E}\left[\widehat{J}_{\mathrm{SIS}}\left({\rho^{\pi}}\right)-\mathbb{E}[{\widehat{J}_{\mathrm{SIS}}\left({\rho^{\pi}}\right)}]-\mathbb{E}[{\frac{1}{N}\sum_{i=1}^{N}\sum_{j=T+1}^{+\infty}\gamma^{j-1}\rho_{1:j}^{\pi}(A_{i,j},S_{i,j})R_{i,j}}]\right]^{2}$
    $=$ $\frac{1}{NT^{2}}\mathbb{V}\left[\sum_{j=1}^{T}\gamma^{j-1}\rho_{1:j}^{\pi}(A_{i,j},S_{i,j})R_{i,j}\right]+\left(\mathbb{E}[{\frac{1}{N}\sum_{i=1}^{N}\sum_{j=T+1}^{+\infty}\gamma^{j-1}\rho_{1:j}^{\pi}(A_{i,j},S_{i,j})R_{i,j}}]\right)^{2}.$
    Denote $\mathcal{H}_{i,t}$ as the $\sigma$-field generated by $\{A_{i,j},S_{i,j},R_{i,j}\}_{j<t}\cup\{A_{i,t},S_{i,t}\}$, and $\mathcal{H}_{i,0}$ as the empty set. A direct calculation yields
    $\mathbb{E}\left(\sum_{j=t}^{+\infty}\gamma^{j-1}\rho_{1:j}^{\pi}(A_{i,j},S_{i,j})R_{i,j}|\mathcal{H}_{i,t}\right)=\gamma^{t-1}\rho_{1:t}^{\pi}(A_{i,t},S_{i,t})Q^{\pi}(A_{i,t},S_{i,t}).$ (D.17)
    Thus, we have
    $\mathbb{E}\left[{\frac{1}{N}\sum_{i=1}^{N}\sum_{j=T+1}^{+\infty}\gamma^{j-1}\rho_{1:j}^{\pi}(A_{i,j},S_{i,j})R_{i,j}}\right]$
    $=$ $\mathbb{E}\left(\mathbb{E}\left[{\sum_{j=T+1}^{+\infty}\gamma^{j-1}\rho_{1:j}^{\pi}(A_{i,j},S_{i,j})R_{i,j}}|\mathcal{H}_{i,T+1}\right]\right)$
    $=$ $\gamma^{T}\mathbb{E}\left(\rho_{1:T+1}^{\pi}(A_{i,T+1},S_{i,T+1})Q^{\pi}(A_{i,T+1},S_{i,T+1})\right)$
    $=$ $\gamma^{T}\mathbb{E}\left[\mathbb{E}\left(\rho_{1:T+1}^{\pi}(A_{i,T+1},S_{i,T+1})Q^{\pi}(A_{i,T+1},S_{i,T+1})\right)|S_{i,T+1}\right]$
    $\leq$ $\gamma^{T}\frac{R_{\max}}{1-\gamma}.$
    Denote $Q_{t}^{\pi,T}(a,s)=\mathbb{E}^{\pi}(\sum_{j=t}^{T}\gamma^{t-1}R_{t}|A_{t}=a,S_{t}=s)$. By the law of total variance in Bowsher and Swain [2012], we have
    $\mathbb{V}\left[\sum_{j=1}^{T}\gamma^{j-1}\rho_{1:j}^{\pi}(A_{i,j},S_{i,j})R_{i,j}\right]$
    $=$ $\sum_{t=1}^{T+1}\mathbb{E}\left[\mathbb{V}\left\{\mathbb{E}\left(\sum_{j=1}^{T}\gamma^{j-1}\rho_{1:j}^{\pi}(A_{i,j},S_{i,j})R_{i,j}|\mathcal{H}_{i,t}\right)|\mathcal{H}_{i,t-1}\right\}\right]$
    $=$ $\sum_{t=1}^{T+1}\mathbb{E}\left[\mathbb{V}\left\{\mathbb{E}\left(\sum_{j=t-1}^{T}\gamma^{j-1}\rho_{1:j}^{\pi}(A_{i,j},S_{i,j})R_{i,j}|\mathcal{H}_{i,t}\right)|\mathcal{H}_{i,t-1}\right\}\right].$ (D.18)
    Since
    $\mathbb{E}\left(\sum_{j=t}^{T}\gamma^{j-1}\rho_{1:j}^{\pi}(A_{i,j},S_{i,j})R_{i,j}|\mathcal{H}_{i,t}\right)=\gamma^{t-1}\rho_{1:t}^{\pi}(A_{i,t},S_{i,t})Q_{t}^{\pi,T}(A_{i,t},S_{i,t}).$ (D.19)
    Combing with (D.19) and (D.2),
    $\mathbb{V}\left[\sum_{j=1}^{T}\gamma^{j-1}\rho_{1:j}^{\pi}(A_{i,j},S_{i,j})R_{i,j}\right]$
    $=$ $\sum_{t=1}^{T+1}\mathbb{E}\left[\mathbb{V}\left\{\left(\rho_{1:t-1}^{\pi}(A_{i,t-1},S_{i,t-1})R_{i,t-1}\gamma^{t-1-1}+\gamma^{t-1}\rho_{1:t}^{\pi}(A_{i,t},S_{i,t})Q_{t}^{\pi,T}(A_{i,t},S_{i,t})\right)|\mathcal{H}_{i,t-1}\right\}\right]$
    $=$ $\sum_{t=1}^{T+1}\mathbb{E}\left[\left\{\gamma^{t-1-1}\rho_{1:t-1}^{\pi}(A_{i,t-1},S_{i,t-1})\right\}^{2}\mathbb{V}\left\{R_{i,t-1}+\gamma\rho_{t}^{\pi}(A_{i,t},S_{i,t})Q_{t}^{\pi,T}(A_{i,t},S_{i,t})|\mathcal{H}_{i,t-1}\right\}\right]$
    $=$ $\sum_{t=1}^{T+1}\mathbb{E}\left[\left\{\gamma^{t-1-1}\rho_{1:t-1}^{\pi}(A_{i,t-1},S_{i,t-1})\right\}^{2}\mathbb{V}\left\{R_{i,t-1}+\gamma\rho_{t}^{\pi}(A_{i,t},S_{i,t})Q_{t}^{\pi,T}(A_{i,t},S_{i,t})|A_{i,t-1},S_{i,t-1}\right\}\right]$
    $=$ $\sum_{t=1}^{T}\mathbb{E}\left[\left\{\gamma^{t-1}\rho_{1:t}^{\pi}(A_{i,t},S_{i,t})\right\}^{2}\mathbb{V}\left\{R_{i,t}+\gamma\rho_{t+1}^{\pi}(A_{i,t+1},S_{i,t+1})Q_{t}^{\pi,T}(A_{i,t+1},S_{i,t+1})|A_{i,t},S_{i,t}\right\}\right].$
    It follows from Assumptions 1 and 2 that $Q_{t}^{\pi,T}(A_{i,t},S_{i,t})$ is upper bounded by $R_{\max}(1-\gamma)^{-1}$. Then,
    $\mathbb{V}\left[R_{i,t}+\gamma\rho_{t+1}^{\pi}(A_{i,t+1},S_{i,t+1})Q_{t}^{\pi,T}(A_{i,t+1},S_{i,t+1})|A_{i,t},S_{i,t}\right]\lesssim\frac{4R_{\max}^{2}}{(1-\gamma)^{2}}.$
    Thus, we have
    $\mathrm{MSE}_{\mathrm{SIS}}\leq\frac{4R_{\max}^{2}}{NT^{2}(1-\gamma)^{2}}\sum_{t=1}^{T}\gamma^{2t-2}\mathbb{E}(\rho_{1:t}^{\pi})^{2}+\frac{\gamma^{2T}R_{\max}^{2}}{(1-\gamma)^{2}}.$
    The above proof relies on the assumption that the behavior action is history-independent. In fact, if the behavior action is history-dependent, the result still holds. We simply need to replace $S_{i,t}$ in the previous formula with the history up to time $t$.
D.3 Proof of Lemma D.2
  • The value of $\kappa_{q}$. Denote $\widehat{Q}^{\pi}(a,s)=\varphi(a,s)^{\top}\widehat{\beta}$ and ${Q}^{\pi}(a,s)=\varphi(a,s)^{\top}\beta^{*}$. Then $\beta^{*}=\Sigma^{-1}\mathbb{E}[R\varphi(A,S)].$ By definition, we have
    $\widehat{\beta}-{\beta}^{*}$
    $=$ $\widehat{{\Sigma}}^{-1}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}{R_{i,j}\varphi(A_{i,j},S_{i,j})}-\widehat{\Sigma}^{-1}\widehat{\Sigma}\beta^{*}$
    $=$ $\widehat{\Sigma}^{-1}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}\varphi(A_{i,j},S_{i,j})\{R_{i,j}-[\varphi(A_{i,j},S_{i,j})-\gamma\sum_{a}\pi(a|S_{i,j+1})\varphi(a,S_{i,j+1})]\beta^{*}\}$
    $=$ $\widehat{\Sigma}^{-1}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}\varphi(A_{i,j},S_{i,j})\{R_{i,j}-[\varphi(A_{i,j},S_{i,j})\beta^{*}-\gamma\sum_{a}\pi(a|S_{i,j+1})\varphi(a,S_{i,j+1})\beta^{*}]\}$
    $=$ $\widehat{\Sigma}^{-1}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}\varphi(A_{i,j},S_{i,j})\{R_{i,j}-Q^{\pi}(A_{i,j},S_{i,j})+\gamma\sum_{a}\pi(a|S_{i,j+1})Q^{\pi}(a,S_{i,j+1})\}$
    $=$ $\underbrace{\left\{\widehat{\Sigma}^{-1}-\Sigma^{-1}\right\}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}\varphi(A_{i,j},S_{i,j})\{R_{i,j}-Q^{\pi}(A_{i,j},S_{i,j})+\gamma\sum_{a}\pi(a|S_{i,j+1})Q^{\pi}(a,S_{i,j+1})\}}_{I_{1}}$
    $+\underbrace{\Sigma^{-1}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}\varphi(A_{i,j},S_{i,j})\{R_{i,j}-Q^{\pi}(A_{i,j},S_{i,j})+\gamma\sum_{a}\pi(a|S_{i,j+1})Q^{\pi}(a,S_{i,j+1})\}}_{I_{2}}.$ (D.20)
    Then
    $\mathbb{E}\|\widehat{\beta}-\beta\|_{2}^{2}\leq\mathbb{E}I_{1}^{2}+\mathbb{E}I_{2}^{2}+2\sqrt{\mathbb{E}I_{2}^{2}\mathbb{E}I_{1}^{2}}.$ (D.21)
    Let
    ${\varepsilon}_{i,j}=\{R_{i,j}-Q^{\pi}(A_{i,j},S_{i,j})+\gamma\sum_{a}\pi(a|S_{i,j+1})Q^{\pi}(a,S_{i,j+1})\}.$
    Let $\mathcal{F}_{i,t}$ denote the sub-$\sigma$-field generated by $\left\{S_{i,t},A_{i,t}\right\}\cup\left\{\left(R_{i,j},A_{i,j},S_{i,j}\right)\right\}_{1\leq j<t}$. Since the data is Markov, we have
    $\mathbb{E}\left(\varepsilon_{i,j}\mid\mathcal{F}_{i,j}\right)=\mathbb{E}\left(\varepsilon_{i,j}\mid S_{i,j},A_{i,j}\right)=0.$
    Notice that $\varphi(A_{i,j},S_{i,j})$ is a function of $S_{i,j}$ and $A_{i,j}$ only, we have for any $1\leq j_{1}<j_{2}\leq T$ that
    $\mathbb{E}[\varepsilon_{i,j_{1}}\varepsilon_{i,j_{2}}\varphi(A_{i,j_{1}},S_{i,j_{2}})^{\top}\varphi(A_{i,j_{2}},S_{i,j_{2}})]=\mathbb{E}\{\varepsilon_{i,j_{1}}\varphi(A_{i,j_{1}},S_{i,j_{1}})^{\top}\varphi(A_{i,j_{2}},S_{i,j_{2}})\mathbb{E}\left(\varepsilon_{i,j_{2}}\mid\mathcal{F}_{i,j_{2}}\right)\}=0.$
    By the independence assumption, we have for any $1\leq j_{1}<j_{2}\leq T$ and $1\leq i_{1}<i_{2}\leq n$ that $\mathbb{E}[\varepsilon_{i_{1},j_{1}}\varepsilon_{i_{2},j_{2}}\varphi(A_{i_{1}},S_{j_{1}})^{\top}\varphi(A_{i_{2}},S_{j_{2}})]=0$. It follows that
    $\mathbb{E}\left\|\sum_{i=1}^{N}\sum_{j=1}^{T}\varphi(A_{i,j},S_{i,j})\varepsilon_{i,j}\right\|_{2}^{2}=\sum_{i=1}^{N}\sum_{j=1}^{T}\mathbb{E}\varepsilon_{i,j}^{2}\varphi_{i,t}^{\top}\varphi_{i,t}=N\sum_{j=1}^{T}\mathbb{E}\varepsilon_{1,j}^{2}\varphi(A_{1,j},S_{1,j})^{\top}\varphi(A_{1,j},S_{1,j}).$
    By Assumptions 1, we obtain
    $\mathbb{E}\left\|\sum_{i=1}^{N}\sum_{j=1}^{T}\varphi(A_{i,j},S_{i,j})\varepsilon_{i,j}\right\|_{2}^{2}\leq(NT)\frac{R_{\max}^{2}}{(1-\gamma)^{2}}\mathbb{E}[\varphi(A_{1,j},S_{1,j})^{\top}\varphi(A_{1,j},S_{1,j})].$ (D.22)
    In the following, we try to bound $\lambda_{\max}(\widehat{\Sigma}^{-1}-\Sigma^{-1})$, where $\lambda_{max}$ denotes the singular values of a matrix. Define ${U}(S)=\sum_{a}\pi(a|S)\varphi(a,S)$, and
    ${\xi}_{i}=\frac{1}{T}\sum_{t=1}^{T}\varphi\left(A_{i,t},S_{i,t}\right)\left\{\varphi\left(A_{i,t},S_{i,t}\right)-\gamma{U}\left(S_{i,t+1}\right)\right\}^{\top}.$
    Denote $L=\max_{a,s}\left\|\varphi\left(a,s\right)\right\|_{2}^{2}$, then $\max_{s}\left\|U\left(s\right)\right\|_{2}^{2}\leq\sqrt{L}$. It follows that
    $\max_{1\leq i\leq N}\lambda_{\max}\left(\xi_{i}-\mathbb{E}\xi_{i}\right)\leq\frac{2}{T}\sum_{t=1}^{T}\sqrt{L}\left(\sqrt{L}+\gamma\sqrt{L}\right)\leq 4L.$ (D.23)
    Let
    $\sigma^{2}$ $=\max\left\{\lambda_{\max}\left(\sum_{i=1}^{N}\mathbb{E}\left(\xi_{i}-\mathbb{E}\xi_{i}\right)\left(\xi_{i}-\mathbb{E}\xi_{i}\right)^{\top}\right),\lambda_{\max}\left(\sum_{i=1}^{N}\mathbb{E}\left(\xi_{i}-\mathbb{E}\xi_{i}\right)^{\top}\left(\xi_{i}-\mathbb{E}\xi_{i}\right)\right)\right\}.$
    By the eigenvalue properties, we can get $\sigma^{2}\leq 16NL^{2}$. Combining this together with (D.23), an application of the matrix concentration inequality (see Theorem 1.6 in Tropp [2012]) yields that
    $\mathbb{P}\left(\lambda_{\max}\left(\frac{1}{N}\sum_{i=1}^{N}(\xi_{i}-\Sigma)\right)\geq\tau\right)\leq 2d\exp\left(-\frac{N^{2}\tau^{2}}{16NL^{2}+8L\tau N/3}\right),\quad\forall\tau>0.$
    Set $\tau=\sqrt{CN\log(N)}L$. For sufficiently large $N$, we have $8L\tau/3\ll\tau^{2}$ and hence
    $\mathbb{P}\left(\lambda_{\max}\left(\sum_{i=1}^{N}\frac{1}{N}(\xi_{i}-\Sigma)\right)\geq\sqrt{CN\log(N)}\frac{L}{N}\right)\leq\frac{2d}{N^{4}}.$
    Since $d\ll N$, we obtain $2d/N^{4}\ll 1/\left(N^{2}\right)$. Thus, we can show that the following event occurs with probability at least $1-O\left(N^{-2}\right)$,
    $\lambda_{\max}(\widehat{\Sigma}-\Sigma)=\lambda_{\max}\frac{1}{N}\sum_{i=1}^{N}\left(\xi_{i}-\Sigma\right)\leq O(\sqrt{N^{-1}\log(N)}L).$ (D.24)
    Now, we have shown $\lambda_{\max}(\widehat{\Sigma}-\Sigma)=O_{p}(\sqrt{N^{-1}\log(N)}L)$. Under the condition that $L=o\{\sqrt{N/\log(N)}\}$, we have $\lambda_{\max}(\widehat{\Sigma}-\Sigma)=o_{p}(1)$. By definition, this implies that $\lambda_{\max}(\widehat{\Sigma}-\Sigma)\leq c/2$, with probability $1-O(N^{-2})$. If follows from Assumption 5, and Cauchy-Schwarz inequality, we have that $\widehat{\Sigma}$ is invertible and satisfies $\lambda_{\max}\left(\widehat{\Sigma}^{-1}\right)\leq 2{c}^{-1}$, with probability $1-O(N^{-2})$. Therefore
    $\lambda_{\max}\left(\widehat{\Sigma}^{-1}-\Sigma^{-1}\right)=$ $\lambda_{\max}\left(\widehat{\Sigma}^{-1}(\widehat{\Sigma}-\Sigma)\Sigma^{-1}\right)$
    $\leq$ $\lambda_{\max}\left(\widehat{\Sigma}^{-1}\right)\lambda_{\max}(\widehat{\Sigma}-\Sigma)\lambda_{\max}\left(\Sigma^{-1}\right)$
    $\leq$ $2{c}^{-2}\lambda_{\max}(\widehat{\Sigma}-\Sigma)$
    with probability $1-O(N^{-2})$. By (D.24), we obtain $\lambda_{\max}\left(\widehat{\Sigma}^{-1}-\Sigma^{-1}\right)\leq O\left\{LN^{-1/2}\log(N)\right\}$ with probability $1-O(N^{-2})$. Combing (D.23), we can get
    $\mathbb{E}[I_{1}^{2}]\leq\{LN^{-1/2}\log(N)+\frac{L}{N^{2}}\}\frac{1}{NT}\frac{R_{\max}^{2}}{(1-\gamma)^{2}}\mathbb{E}[\varphi(A_{1,j},S_{1,j})^{\top}\varphi(A_{1,j},S_{1,j})].$ (D.25)
    It follows from (D.22) and Assumption (5) that
    $\mathbb{E}[I_{2}^{2}]\leq c^{-2}\frac{1}{NT}\frac{R_{\max}^{2}}{(1-\gamma)^{2}}\mathbb{E}[\varphi(A_{1,j},S_{1,j})^{\top}\varphi(A_{1,j},S_{1,j})].$ (D.26)
    Combing (D.21), (D.25) and (D.26) that
    $\mathbb{E}\|\widehat{\beta}-\beta\|_{2}^{2}\leq\frac{1}{NT}\frac{R_{\max}^{2}}{(1-\gamma)^{2}}\mathbb{E}[\varphi(A_{1,j},S_{1,j})^{\top}\varphi(A_{1,j},S_{1,j})].$ (D.27)
    It follows from the definition of $Q^{\pi}$ and $\widehat{Q}^{\pi}$, we can get
    $\mathbb{E}\left\{(\widehat{Q}^{\pi}(a,s)-Q^{\pi}(a,s))^{2}\right\}$ $\leq\varphi(a,s)^{\top}\varphi(a,s)\mathbb{E}\lambda_{\max}[(\widehat{\beta}-\beta)(\widehat{\beta}-\beta)^{\top}]$
    $=$ $\varphi(a,s)^{\top}\varphi(a,s)\frac{1}{NT}\frac{R_{\max}^{2}}{(1-\gamma)^{2}}\mathbb{E}[\varphi(A_{1,j},S_{1,j})^{\top}\varphi(A_{1,j},S_{1,j})].$ (D.28)
    Thus, we can get
    $\mathbb{E}_{A,S}\mathbb{E}\left\{(\widehat{Q}^{\pi}(A,S)-Q^{\pi}(A,S))^{2}\right\}$
    $\leq$ $\mathbb{E}_{A,S}[\varphi(A,S)^{\top}\varphi(A,S)]\frac{1}{NT}\frac{R_{\max}^{2}}{(1-\gamma)^{2}}\mathbb{E}[\varphi(A_{1,j},S_{1,j})^{\top}\varphi(A_{1,j},S_{1,j})].$
    It follows from the definition of $\kappa_{q}$, we have
    $\kappa_{q}=\sqrt{\frac{1}{NT}\mathbb{E}_{A,S}[\varphi(A,S)^{\top}\varphi(A,S)]\mathbb{E}[\varphi(A_{1,j},S_{1,j})^{\top}\varphi(A_{1,j},S_{1,j})]}.$ (D.29)
    In tabular mode, we have $\mathbb{E}_{A,S}[\varphi(A,S)^{\top}\varphi(A,S)]\leq d=|\mathcal{S}||\mathcal{A}|$, where ||$\cdot$|| ⋅ | denotes the cardinality of a set. Then, we have
    $\kappa_{q}=\sqrt{\frac{1}{NT}}|\mathcal{S}||\mathcal{A}|.$ (D.30)
  • The value of $\kappa_{w}$.
    Denote ${w}^{\pi}(a,s)=\varphi(a,s)^{\top}\alpha^{*}$ with $\alpha^{*}$ satisfying $\alpha^{*}=\Sigma^{-1}(1-\gamma)\mathbb{E}[\sum_{a}\pi(a|S_{i,1})\varphi(a,S_{i,1})].$ By definition, we have
    $\widehat{\alpha}-{\alpha}^{*}$
    $=$ $\widehat{{\Sigma}}^{-1}(1-\gamma)\frac{1}{N}\sum_{i=1}^{N}\left[\sum_{a}\pi(a|S_{i,1})\varphi\left(a,S_{i,1}\right)\right]-\widehat{\Sigma}^{-1}\widehat{\Sigma}\alpha^{*}$
    $=$ $\widehat{\Sigma}^{-1}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}\left\{(1-\gamma)\sum_{a}\pi(a|S_{i,1})\varphi(a,S_{i,1})-w^{\pi}(A_{i,j},S_{i,j})\right.$
    ×$\left.\times[\varphi(A_{i,j},S_{i,j})-\gamma\sum_{a}\pi(a|S_{i,j+1})\varphi(a,S_{i,j+1})]\right\}$
    $=$ $\left\{\widehat{\Sigma}^{-1}-\Sigma^{-1}\right\}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}L_{i,j}+\Sigma^{-1}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}L_{i,j},$ (D.31)
    where
    $L_{i,j}=$$\left\{(1-\gamma)\sum_{a}\pi(a|S_{i,1})\varphi(a,S_{i,1})-w^{\pi}(A_{i,j},S_{i,j})\right.$
    ×$\left.\times[\varphi(A_{i,j},S_{i,j})-\gamma\sum_{a}\pi(a|S_{i,j+1})\varphi(a,S_{i,j+1})]\right\}.$
    Then, direct calculations yield
    $\mathbb{V}[\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}L_{i,j}]=$ $\frac{1}{N}\mathbb{V}[\frac{1}{T}\sum_{j=1}^{T}L_{i,j}]$
    $=$ $\frac{1}{N}\mathbb{V}\left[(1-\gamma)\sum_{a}\pi(a|S_{i,1})\varphi(a,S_{i,1})-\frac{1}{T}\sum_{j=1}^{T}w^{\pi}(A_{i,j},S_{i,j})\right.$
    ×$\left.\times[\varphi(A_{i,j},S_{i,j})-\gamma\sum_{a}\pi(a|S_{i,j+1})\varphi(a,S_{i,j+1})]\right].$
    Then by similar arguments as (D.29), we can get
    $\kappa_{w}=\sqrt{\frac{1}{N}\mathbb{E}_{A,S}\varphi(A,S)^{\top}\varphi(A,S)\mathbb{E}[\varphi(A_{i,j},S_{i,j})^{\top}\varphi(A_{i,j},S_{i,j})]}.$ (D.32)
    In tabular mode, we have
    $\kappa_{w}=\sqrt{\frac{1}{N}}|\mathcal{S}||\mathcal{A}|.$ (D.33)